# Noncollinear Spin-Flip TDDFT for Potential Energy Surface Crossings: Conical Intersections and Spin Crossings

### v2

@@ 4 Assessment of the Potential of Noncollinear Spin-Flip TDDFT for Nonadiabatic Dynamics Simulations
    Inclusion of doubly-excited configuration $|{\mathrm{DES}}\rangle$. Finally, we specifically highlight the treatment of $|{\mathrm{DES}}\rangle$, as these are notoriously difficult to describe within spin-conserving TDDFT89. The inclusion of $|{\mathrm{DES}}\rangle$ actually plays an important role in shaping the topology of $\mathrm{S}_{0}/\mathrm{S}_{1}$ conical intersections, preventing spurious PES crossings29, and explaining inverted singlet-triplet gaps72. Fortunately, noncollinear spin-flip TDDFT naturally includes $|{\mathrm{DES}}\rangle$. A direct way to assess the impact of $|{\mathrm{DES}}\rangle$ is to examine electronic states dominated by it. Four molecular systems, as shown in Figure 7, are selected here for illustrative purposes. The geometry is optimized at the RKS B3LYP/6-31G* level in Gaussian 16, and the vertical excitation energies of states dominated by doubly excited character are computed using noncollinear spin-flip TDDFT. The results in Table 4 show reasonable agreement with the reference values, indicating that noncollinear spin-flip TDDFT can provide a reliable description of states dominated by $|{\mathrm{DES}}\rangle$. It is worth noting that in DHQQ-Et, the partition of $|{\mathrm{DES}}\rangle$ is relatively low and features a significant multiconfigurational character, making it challenging to describe.
    Figure 7: Chemical structures of the four test systems. (DHQQ-Et stands for 5,12-diethyl-5,12-dihydroquinoxalino[2,3-b]quinoxaline.)
@@
    Table 4: Excitation energies of states dominated by doubly excited character and partition of $|{\mathrm{DES}}\rangle$ configuration calculated by noncollinear spin-flip TDDFT.
@@
    Nitrous acid 7.90 7.97a 0.90
    Cyclobutadiene 5.08 4.04a 0.95
    Oxalyl Fluoride 9.58 8.92a 0.97
    DHQQ-Et 5.80 2.58b 0.62
    aTheoretically best estimate values from Ref. 90. bExperimental value from Ref. 91.
  8. 8.
    Computational efficiency. Computational cost is always a key consideration for practical dynamics simulations. As discussed in Ref. 50, the computational cost of noncollinear spin-flip TDDFT and its gradients is comparable to that of spin-conserving TDDFT, significantly lower than that of traditional multireference wavefunction methods.
Overall, noncollinear spin-flip TDDFT is a remarkably clean and simple approach. Without introducing any ad hoc treatments, it can correctly capture the topological features near potential energy surface crossings and provide reasonable spin multiplicities. Moreover, these features are not dependent on the choice of a specific functional. Although the method still faces issues, such as slight spin contamination in low-lying excited states and inadequate spin description for some high-lying states, its low computational cost still makes it a promising electronic structure tool for nonadiabatic dynamics simulations, especially for larger systems.
@@
  • Cartesian coordinates of: conical intersections of ethylene, PSB3, and styrene; $\mathrm{S}_{1}/\mathrm{T}$ crossing point of $o$-nitrophenol; and the optimized structures of cyclazine, heptazine, nitrous acid, cyclobutadiene, oxalyl fluoride, and DHQQ-Et obtained in this work (PDF)

### v3

@@ Appendix B Calculations of Excited States Dominated by Doubly Excited Configuration
The $|{\mathrm{DES}}\rangle$ configuration plays important role in shaping the topological features of $\mathrm{S}_{0}/\mathrm{S}_{1}$ conical intersectionsteh2019simplest, avoiding artificial potential energy surface crossingsteh2019simplest, and explaining inverted singlet-triplet gapsdesilva2019inverted. Although it is missed by conventional spin-conserving TDDFTmaitra2004double, it is naturally included in noncollinear spin-flip TDDFT as a singly-excited state with respect to the reference state. A direct way to assess the impact of $|{\mathrm{DES}}\rangle$ is to examine electronic states dominated by it. Three molecular systems, as shown in Figure 8, are selected here for illustrative purposes. The geometry is optimized at the B3LYP/cc-pVTZ level in Gaussian 16g16 and the vertical excitation energies of states dominated by doubly excited character are computed using noncollinear spin-flip TDDFT. The results in Table 6 show reasonable agreement with the reference values, indicating that noncollinear spin-flip TDDFT can provide a reliable description of states dominated by $|{\mathrm{DES}}\rangle$.
Figure 8: Chemical structures of the three test systems.
@@
Table 6: Excitation energies (eV) of states dominated by doubly excited character and partition of $|{\mathrm{DES}}\rangle$ configuration calculated by noncollinear spin-flip TDDFT.
@@
Nitrous acid 7.95 7.97a 0.89
Cyclobutadiene 4.97 4.04a 0.95
Oxalyl Fluoride 9.68 8.92a 0.97
aTheoretically best estimate values from Ref. kossoski2024reference.
Appendix C Single Bond Dissociation
[…]
@@ Appendix C Single Bond Dissociation
  • Cartesian coordinates of MECIs and MECPs located in this work.
  • Cartesian coordinates of geometries used in Table 6.
  • Computational cost comparison between noncollinear spin-flip TDDFT and collinear spin-flip TDDFT.